OpenStudy (anonymous):

hi!!!!! how to solve 5x - 4 = 9x + 3 ??? <3

OpenStudy (anonymous):

just a number as answer please

OpenStudy (anonymous):

-7/4

OpenStudy (anonymous):

5x -4 = 9x+3 5x-9x=3+4 -4x = 7 x = -7/4